OpenStudy (anonymous):

If f(x) = x - 1 and g(x) = 2x - 3, find (f + g)(x). 6. If f(x) = x - 1 and g(x) = 2x - 3, find (fg)(x). 7. If f(x) = x - 1 and g(x) = 2x - 3, find (g - f)(x). 8. If f(x) = x + 4 and g(x) = 5x - 2, find (g-f)(2).

OpenStudy (anonymous):

ill give a medal to helper

OpenStudy (anonymous):

hey

OpenStudy (mathmale):

"(f + g)(x)" represents the SUM of two functions f and g. Line them up: f(x) = x - 1 g(x) = 2x - 3 Here you see 3 columns. Add each column up. Write the sum of the first column as the label to your answer: (f+g)(x) = ? - ?

OpenStudy (anonymous):

2^x2 -1-3?